E is 54 more than F but E+F=90 find the measure of E and F
Answer by greenestamps(13200)



(1) E = F+54 (E is 54 more than F)

(2) E+F = 90

Use (1) to substitute "F+54" for "E" in (2):

(F+54)+F = 90

Solve using basic algebra.